OpenStudy (anonymous):

Write the equation of the line with slope 5/3 which passes through the point (3, -2).

OpenStudy (anonymous):

Yokieee can u answer on my question :P

OpenStudy (anonymous):

y-y1=m(x-x1), in this case m=5/3 and (x1,y1)=(3,-2)

OpenStudy (anonymous):

you have to plug the numbers into the formula

OpenStudy (anonymous):

totally wrong :P

OpenStudy (anonymous):

it should look like this: y-(-2)=5/3(x-3), now just evluate that

OpenStudy (anonymous):

is this right y = (5/3)x + 3

OpenStudy (anonymous):

you got the slope right

OpenStudy (anonymous):

y + 2 = 5/3x +3

OpenStudy (anonymous):

but check the arithmetic

OpenStudy (anonymous):

wait..did that wrong lol

OpenStudy (anonymous):

y=5/3x-7

OpenStudy (anonymous):

y + 2 = 5/3(x - 3)

OpenStudy (anonymous):

y + 2 = 5/3x -5

OpenStudy (anonymous):

y = 5/3x - 7

OpenStudy (anonymous):

there we go

OpenStudy (anonymous):

yup i got the same answer y=5/3x-7